OpenStudy (danysanchy97):

help please The product of 8/15, 6/5, 1/3 is A. 48/15. B. 48/30. C. 16/75. D. 16/15.

OpenStudy (danysanchy97):

@mathmale @pooja195

OpenStudy (mathmale):

What does "product" mean? Please write the product "of 8/15, 6/5, and 1/3" without the word "and."

OpenStudy (mathmale):

Hint: the product of 16 and 3 is 16(3)=48, or (16)(3)=48

OpenStudy (mathmale):

Please show the work you did that makes you think the answer is C.

OpenStudy (danysanchy97):

can you walk me through the answer? i'm not that good with fractions. that was just a guess

OpenStudy (danysanchy97):

i just want help on this question

OpenStudy (mathmale):

I'll make a deal with you. You do some work, post some results, and then I'll help you. Please stop your "I'm not that good" comments; it helps no one, least of all you.

OpenStudy (mathmale):

What does the word "product" mean in this context?

OpenStudy (danysanchy97):

It means multiply.

OpenStudy (mathmale):

Write "product of 8/15, 6/5, 1/3" symbolically (no commas). "product of 8/15, 6/5, 1/3" = ??

OpenStudy (mathmale):

Go ahead and multiply to find product of 8/15, 6/5, 1/3.

OpenStudy (mathmale):

Again: without commas.

OpenStudy (danysanchy97):

\[\frac{ 8 }{ 15 } \times \frac{ 6 }{ 5 } \times\frac{ 1 }{ 3 }\]

OpenStudy (danysanchy97):

is that how i set it up?

OpenStudy (mathmale):

Yes! Excellent! Now, look for ways to factor the numerator and denominator and to reduce the fraction BEFORE trying to multiply. For example, 18=2*3*3

OpenStudy (mathmale):

This is a product, so discard any consideration of LCD (Lowest Common Denominator).

OpenStudy (mathmale):

Hint: do 15 and 6 have any common factor or factors?

OpenStudy (danysanchy97):

\[\frac{48 }{ ?225 }= \frac{ 16 }{ 75 }\]

OpenStudy (danysanchy97):

is it correct? i didnt mean to put the question mark on there

OpenStudy (mathmale):

You are multiplying here. While that's OK, I had previously asked you to use a simpler method: FACTOR all numerators and denominators in the given fractions and then cancel wherever possible. Your approach is OK, but mine is shorter and simpler.

OpenStudy (danysanchy97):

oh ok. but is the answer correct?